OpenStudy (anonymous):

Solve for x? x^2/3 – 2x^1/3 – 8 = 0

OpenStudy (anonymous):

let u=x^{1/2}

OpenStudy (anonymous):

if you do this what do you get?

OpenStudy (anonymous):

Hmm.. I seem to be having trouble with this still. I'm not quite sure how to set it up. I'm pretty lost with it. :P Thanks for helping.

OpenStudy (anonymous):

actually i mean't if you let u=x^{1/3}

OpenStudy (anonymous):

if you do that you get \[u^2-2u-8=0\]

OpenStudy (anonymous):

Sooo... it told me to solve for x, so how exactly did the u come in? I'm sorry. My brain is kind of shutting down right now after so many hours of math work. :P

OpenStudy (anonymous):

u is just a variable to simplify the problem

OpenStudy (anonymous):

Oh, okay. :) That makes sense.

OpenStudy (anonymous):

so now you have \[u^2-2u-8\] can you factor this?

OpenStudy (anonymous):

Yes. Is this right? (u-4)(u+2)

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

so now your roots are 4 and -2 (from the equations u-4=0 and u+2=0)

OpenStudy (anonymous):

however these are the roots for u so now you have to plug these into \[u=x^{1/3}\]

OpenStudy (anonymous):

you get \[4=x^{1/3}\] \[-2=x^{1/3}\]

OpenStudy (anonymous):

Ohh.... okay. :) So what would I go about doing at this point then?

OpenStudy (anonymous):

cube both sides to get x by itself

OpenStudy (anonymous):

\[4^3=(x^{1/3})^3\]

OpenStudy (anonymous):

\[64=x\]

OpenStudy (anonymous):

now for the next one you have \[-2^3=(x^{1/3})^3\]

OpenStudy (anonymous):

\[-8=x\]

OpenStudy (anonymous):

so your solution set is [64,0],[-8,0]

OpenStudy (anonymous):

Awesome. :) I really appreciate your help on this SOOOOO much. Thank you, thank you, thank you.

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

So if I close this question will you still be given a medal? This is my first time asking a question and I certainly want you to be rewarded for your help. :)

OpenStudy (anonymous):

yes you can close if you like